A computer repair service charges $50 to come out for a service call which is applied toward the first 30 minutes of service should a repair be needed. If a repair is needed, the charge is $37 per 30 minutes (after the initial half hour included in the service call charge)for labor and the cost of parts. Write a cost function to determine the amount a customer needing a repair that takes 1.5 hours and requires $103 part will be charged.
Answer by checkley71(8403)

COST=FIRST HALF HOUR SERVICE CHARGE(50)+ NUMBER OF 30 MINUTE PERIODS(2) TIMES THE RATE (37)+ THE PRICE OF THE PARTS (103)
COST=50+2*37+103=$227

FIRST HALF HOUR=50
NEXT HOUR=37*2=74
PARTS=103
50+37*2+103=50+74+103=$227 TOTAL COST FOR SERVICE CALL.